Nordic Optical Telescope

Search
Text Size:
Smaller Text Normal Text Larger Text

Facebook

Follow us on Twitter



next up previous contents
Next: Computers Up: AiC Report to NOT Previous: Telescope

Instruments

ALFOSC

Grism wheel timeouts

ALFOSC has performed well with the exception of the occasional timeout from the grism wheel. Normally this is easily solved by stopping the user interface and initializing on restart, occasionally the telescope must be brought to zenith first. The final cause of these timeouts is not clear due to their sparsely intermittent nature, but is believed to be linked to unevenly distributed weight when only partially populated, and appears to depend on telescope altitude. Close inspection has yielded no anomalies.

Mounting

CUO/IJAF noted that the ALFOSC flange did not fit snugly to FASU due to the absence of a mounting bolt at the detector end of the instrument. The resulting lever arm made it impossible to completely close the gap between the flange and FASU. This has been a known problem at the telescope for some time and was not believed to be a significant issue. However, CUO/IJAF were concerned that the FASU flange would be permanently bent so a fitting was produced to allow the inclusion of a mounting bolt near the detector. The flange and FASU are now flush when the instrument is properly mounted and no permanent damage can be discerned (CP, PB).

Cleaning

During January technical time ALFOSC was thoroughly cleaned and inspected. A considerable amount of dust was removed from the interior and most particularly from the upper surface of the folding prism. The resulting flat field shows considerable improvement - where there were as many as one hundred dust donuts, immediately after the cleaning there were less than 10. Complete removal of dust was not possible without dismounting the CCD, which was not considered necessary.

MOS Mode

In collaboration with Per Kjaergaard (CUO), the MOS mode is now nearing the condition at which it may be made available to the community. We expect to announce this in the call for proposals made in April and support runs in the semester beginning October 2001. Remaining work is primarily in the area of documentation. Runs will require significantly more support than usual, and we recommend a restrained response to demand in the first instance. (JC)

Zero-Point Monitoring Programme

Data has been collected over several months and is being reduced by Andreas Jaunsen (ESO) as part of the gravitational lens monitoring programme led by Jen Hjorth (CUO). Some results are already posted on the NOT web pages. (AK)

TurPol

New ND filters installed and calibrated (VP). New document set available on web (AK). Data reduction programmes now available on web page (AK). The ``circular polarimetry only'' mode has been tested and reduction software written by A. Berdyugin is available.

HiRAC

There is no change in the disposition of HiRAC - observers are being requested to use ALFOSC where possible and so far this has always been the case. No attempt has been made to rectify the soaking problem due to lack of manpower and the anticipated imminent arrival of MOSCA which will make HiRAC redundant.

StanCam

StanCam continues to operate without problems.

FIES

A new fibre-viewing mirror and mount have been installed in the telescope instrument adapter, including new StanCam reimaging optics to aid target acquisition. A new fibre mounting plate has been installed. Target acquisition procedures have been established and test observations made under real observing conditions.

Now FIES is ready for final commissioning: All hardware is installed. Target acquisition procedures have been described and tested, and can be easily carried out by the observer.

FIES still needs a new CCD, as the test-chip proved to have very high dark current. The mounting of a new CCD will not interfere with normal telescope operation. On the software side a smoother interface to BIAS would be desirable but not critical. Also the instrument needs to be put back on the telescope pier. Apart from this, FIES is ready for use. (TD & FIES team)

SOFIN

One SOFIN CCD suffered mechanical destruction and both have been replaced with a similar devices donated by Isaac Newton Group. (GC)

NOTCam

The mechanically destroyed engineering grade device has been replaced. Before it is installed, the CUO/IJAF group are investigating detector mount redesigns so as to avoid repeating the circumstances which destroyed the first. A delivery date to La Palma is unknown at this time.

MOSCA

Five (U', B, V, R and SDSS z') of the interim filter set have been received and test data collected in parallel with the ALFOSC zero point monitoring programme.

FRed

See Director's report.

CCD Acquisition

In line with the working group's recommended priorities, we have given priority to obtaining state-of-the-art detectors first for MOSCA and next for ALFOSC. The best candidates for MOSCA are currently Marconi (ex-EEV) devices with integrated anti-fringing technology, but the possibility of a good device (4k tex2html_wrap_inline396 4k monolithic) and price from Fairchild prevent us from committing to Marconi. Fairchild could also provide a replacement detector for ALFOSC.

All investigations of best-effort, wafer-run purchases have been postponed in the apparent absence of a suitable path of manufacture, packaging and test. Moreover, we are not aware of any suitable partners available to share the expense and risk at this time.

High speed CCD controller serial link

See detailed status report from GC in Appendix B.

Progress is being made. A visiting student, Bjarne Andersen, will assist with the project from mid-March to mid-June. It is hoped that CUO/IJAF will integrate this design into their planned improvements to their system.


next up previous contents
Next: Computers Up: AiC Report to NOT Previous: Telescope

Tim Abbott, AiC
Wed Mar 21 12:56:05 GMT 2001

Back to top Last modified: 21-Mar-2001